Chapter 31, Problem 045 (a) In an RLC circuit, can the amplitude of the voltage across an inductor be greater than the amplitude of the generator emf? (b) Consider an RLC circuit with driving emf amplitude Em-8 V resistance R = 9 Ω, inductance L = 1.0 H, and capacitance C = 1.1 μF. Find the amplitude of the voltage across the inductor at resonance (b) Number Units the tolerance is +/-596
Answer #1

a) Yes

the statement in the first question it is right. You can check in Answer of the second question.

b)

At resonance

Angular frequency   w^2= 1/LC

here L=1.0 H , C=1.1 ×10^(-6) F

w^2=90.91×10^4

w=953.4 7 rad/s

at resonance current in the circuit

I=Em/R =8/9

I= 0.89 A

voltage across inductor

V =I×X here X=wL

V=0.89×953.47×1

V=848.5883

V=848.59 V. Answer